# references.bib

### v1

@@ 3.3 Overall Performance and Model Comparison
Figure 1: Average CTRS Category Scores by Model Variant
@@ 3.4 Mixed Effects Model Analysis
C(model)[T.Mistral] -8.054 1.050 -7.667 $0$10.113 -5.995
C(model)[T.Qwen] -4.160 1.041 -3.997 $0$6.200 -2.120
C(variant)[T.instruct] -11.326 0.855 -13.246 $0$13.002 -9.650
@@ 3.5 Consistency
Figure 2: Distribution of Total CTRS Scores by Model Variant
@@ 3.6 CTRS Category Analysis
Analysis of individual CTRS categories revealed distinct patterns of strengths and weaknesses across both model types and variants (CBT-tuned vs instruct). Figure 4 illustrates the distribution of scores for each model within the CTRS domains. CBT-tuned Llama 3.1 8b was strong across all categories but scored particularly high on Understanding (6.83 ± 0.38), Interpersonal Effectiveness (6.93 ± 0.29), and Focusing on Key Cognitions or Behaviors (6.87 ± 0.49).
@@
Figure 3: Distribution of Scores by Model Variant and CTRS Domain
4.   Discussion
@@ 4.2 Implementation of CBT Techniques
The CBT-tuned models demonstrated reasonable competence in socializing the simulated patients to the CBT model and implementing core CBT techniques such as thought records, Socratic questioning and cognitive distortions. Homework assignments were generally appropriate and aligned with SMART goal principles. The CBT-tuned models also consistently reviewed the previous session’s homework at the start of the current session. However, we observed a tendency to introduce cognitive distortions and cognitive restructuring earlier in the therapy course than might be clinically appropriate [Kazantzis2018TheMeta-Analyses, Fenn2013TheTherapy, McGinn2000CognitiveStatus, Wenzel2017BasicTherapy]. The models also failed to engage in a robust exploration of the patient’s depression symptoms and social history in the initial phase of the therapy [McGinn2000CognitiveStatus]. These limitations likely reflect biases in our training data distribution. One particular strength of the CBT-tuned models, was their consistent acknowledgement of AI limitations at the start of every therapy session. Unfortunately, near the termination of sessions, there were lapses in the models’ abilities to recognize their limitations, as evidenced by offering call and text support to the patient and performing “online” searches for resources. These mistakes suggest significant room for improvement in constraint enforcement.
@@ 4.3 Therapeutic Relationship
Both the instruct-tuned and CBT-tuned models, scored well in the CTRS categories of Interpersonal Effectiveness and Understanding. All model variants made use of empathetic statements, however, the CBT-tuned models were more likely to engage in regular check-ins and collaborative exploration of patient concerns. Despite, demonstrating effective interpersonal skills, the CBT-tuned models often conducted only a superficial exploration of patient concerns, missing opportunities for deeper examination of thoughts and emotions [Kazantzis2018TheMeta-Analyses, Fenn2013TheTherapy].
@@ 4.7 Ethics and Future Directions
Any work on AI based therapy systems necessitates a discussion of the ethical implications [Mirzaei2024ClinicianImplications, Ma2024NoPsycho-counseling]. Our study was meant to demonstrate the technical feasibility and limitations of creating an AI therapist to deliver CBT. We do not recommend deploying any of the models from our study in clinical applications. This recommendation is made based on the significant limitations of these models including, hallucinations, constraint enforcement and long context coherence. We emphasize that these models should be viewed as tools for further research and as a stepping stone for future efforts to create effective AI therapists rather than clinical aids.
Derivative studies may want to use higher quality synthetic data with more rigorous validation procedures, larger models and clinical validation. It should be recognized that that therapy, of any modality, has the potential to be harmful for a patient if delivered incorrectly [Castonguay2010TrainingTreatments., Knox2019TheFails]. As such, clinical validation studies are essential to assess real-world applicability. These studies must carefully consider patient safety and include robust monitoring systems for adverse events. As LLMs and AI therapists advance, we will also need to develop clear guidelines for appropriate use cases and contraindications [Stade2024LargeEvaluation, Obradovich2024OpportunitiesPsychiatry].
Hallucinations remain a fundamental limitation of LLMs. Furthermore, due to their auto-regressive nature and probabilistic sampling of output tokens, their responses may not always align with expectation or clinical indication [Lappin2024AssessingModels, Xu2024HallucinationModels]. Reinforcement learning with human feedback has been very effective in constraining model responses and providing generally appropriate guardrails for content [Lin2024ReinforcementGiants]. However, in sensitive clinical applications with vulnerable patients, even these guardrails may prove insufficient [Roose2024CanSuicide].
5.   Conclusion
This study represents a significant step forward in developing accessible, efficient AI systems capable of delivering CBT-informed interventions. Our results demonstrate that small open-weight language models can be successfully fine-tuned to provide structured therapeutic interactions with reasonable fidelity to CBT principles. The significant improvements observed across all models after CBT fine-tuning suggest that targeted training can effectively encode therapeutic competencies even in models with limited parameters. Future studies should focus on addressing the identified technical limitations, enhancing quality of training data, refining evaluation methodologies, and clinical validation.
@@ 5.   Conclusion
The data, codebase, and implementation details for this study are publicly available under MIT License at GitHub Repository. All fine-tuned models developed in this study are publicly available through the Hugging Face model repository (https://huggingface.co/TTahir).
[…]

### v2

@@ 3.2 Overall performance and model comparison
Figure 1: Average CTRS category scores by model variant. The radar chart visualizes the mean performance score for each model variant across the $1$ Cognitive Therapy Rating Scale (CTRS) categories. Each axis represents a distinct CTRS category, and the distance from the center indicates the average score. Each colored polygon corresponds to a specific model variant.
@@ 3.3 Mixed effects model analysis
C(model)[T.Mistral] -8.054 1.050 -7.667 0.000 -10.113 -5.995
C(model)[T.Qwen] -4.160 1.041 -3.997 0.000 -6.200 -2.120
C(variant)[T.instruct] -11.326 0.855 -13.246 0.000 -13.002 -9.650
@@ 3.4 Consistency
Figure 2: Distribution of total CTRS scores by model variant. The box plot illustrates the distribution of total CTRS scores for each model variant. The central line in each box represents the median score, while the box boundaries indicate the interquartile range (25th to 75th percentiles). Whiskers extend to show the range of the data, and individual points represent outliers.
@@ 3.5 CTRS category analysis
Analysis of individual CTRS categories revealed distinct patterns of strengths and weaknesses across both model types and variants (CBT-tuned vs instruct). CBT-tuned Llama 3.1 8b was strong across all categories but scored particularly high on Understanding (6.83 ± 0.38), Interpersonal Effectiveness (6.93 ± 0.29), and Focusing on Key Cognitions or Behaviors (6.87 ± 0.49).
@@
4 Discussion
@@ 4.2 Implementation of CBT techniques
The CBT-tuned models demonstrated reasonable competence in socializing the simulated patients to the CBT model and implementing core CBT techniques such as thought records, Socratic questioning and cognitive distortions. Homework assignments were generally appropriate and aligned with SMART goal principles. The CBT-tuned models also consistently reviewed the previous session’s homework at the start of the current session. However, we observed a tendency to introduce cognitive distortions and cognitive restructuring earlier in the therapy course than might be clinically appropriate [31, 32, 33, 48]. The models also failed to engage in a robust exploration of the patient’s depression symptoms and social history in the initial phase of the therapy [33]. These limitations likely reflect biases in our training data distribution. One particular strength of the CBT-tuned models, was their consistent acknowledgement of AI limitations at the start of every therapy session. Unfortunately, near the termination of sessions, there were lapses in the models’ abilities to recognize their limitations, as evidenced by offering call and text support to the patient and performing “online” searches for resources. These mistakes suggest significant room for improvement in constraint enforcement.
@@ 4.3 Therapeutic relationship
Both the instruct-tuned and CBT-tuned models, scored well in the CTRS categories of Interpersonal Effectiveness and Understanding. All model variants made use of empathetic statements, however, the CBT-tuned models were more likely to engage in regular check-ins and collaborative exploration of patient concerns. Despite, demonstrating effective interpersonal skills, the CBT-tuned models often conducted only a superficial exploration of patient concerns, missing opportunities for deeper examination of thoughts and emotions [31, 32].
@@ 4.7 Ethics and future directions
Any work on AI based therapy systems necessitates a discussion of the ethical implications [49, 50]. Our study was meant to demonstrate the technical feasibility and limitations of creating an AI therapist to deliver CBT. We do not recommend deploying any of the models from our study in clinical applications. This recommendation is made based on the significant limitations of these models including, hallucinations, constraint enforcement and long context coherence. We emphasize that these models should be viewed as tools for further research and as a stepping stone for future efforts to create effective AI therapists rather than clinical aids.
Derivative studies may want to use higher quality synthetic data with more rigorous validation procedures, larger models and clinical validation. It should be recognized that that therapy, of any modality, has the potential to be harmful for a patient if delivered incorrectly [51, 52]. As such, clinical validation studies are essential to assess real-world applicability. These studies must carefully consider patient safety and include robust monitoring systems for adverse events. As LLMs and AI therapists advance, we will also need to develop clear guidelines for appropriate use cases and contraindications [11, 10].
Hallucinations remain a fundamental limitation of LLMs. Furthermore, due to their auto-regressive nature and probabilistic sampling of output tokens, their responses may not always align with expectation or clinical indication [12, 53]. Reinforcement learning with human feedback has been very effective in constraining model responses and providing generally appropriate guardrails for content [54]. However, in sensitive clinical applications with vulnerable patients, even these guardrails may prove insufficient [55].
5 Conclusion
This study represents a significant step forward in developing accessible, efficient AI systems capable of delivering CBT-informed interventions. Our results demonstrate that small open-weight language models can be successfully fine-tuned to provide structured therapeutic interactions with reasonable fidelity to CBT principles in simulated environments. The significant improvements observed across all models after CBT fine-tuning suggest that targeted training can effectively encode therapeutic competencies even in models with limited parameters. Future studies should focus on addressing the identified technical limitations, enhancing quality of training data, refining evaluation methodologies, and clinical validation.
@@ 5 Conclusion
The data, codebase, and implementation details for this study are publicly available under MIT License in the project’s GitHub Repository. All fine-tuned models developed in this study are publicly available through the Hugging Face model repository - HuggingFace/TTahir.